OpenStudy (anonymous):

what is the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (anonymous):

Well, slope intercept form is y = mx+b So, if you get y alone you can see the slope. 5x-3y=-6 -5x -5x -3y = -5x-6 /-3 /-3 y=-5/3x+2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

remember that -5x / -3 is 5/3 x y = 5/3 x + 2

OpenStudy (anonymous):

Oh yes, thank you @binarymimic good catch!! :o)